OpenStudy (anonymous):

Liam owns a business earning $43,200 in profits. These profits are appreciating at about 1.2% each year. What are Liam’s profits after six years?

OpenStudy (anonymous):

SOMEONE PLEASE HELP ME PLEASE I BEG YOU PLEASE PLEASE!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

First year, 43200 + (1.2/100 x 43200) = 43718.40 Profit = 43718.40 - 43200 Second year, 43718.40 + (1.2/100 x 43718.4) = 44243.02 Profit = 44243.02 - 43718.40 Third year, 44243.02 + (1.2/100 x 44243.02) = 44773.94 Profit = 44773.94 - 44243.02 Fourth year, 44773.94 + (1.2/100 x 44773.94) = 45311.23 Profit = 45311.23 - 44773.94 Fifth year, 45311.23 + (1.2/100 x 45311.23) = 45854.96 Profit = 45854.96 - 45311.23 Sixth year, 45854.96 + (1.2/100 x 45854.96) = 46405.22 Profit = 46405.22 - 45854.96